Acme claims that it takes a mean of 150 hours to produce a shipment of cartons with a standard deviation of 70 hours. We want to col- lect enough data to show that the standard deviation is actually less. In a sample of 300 shipments, the mean time to produce a shipment is 158 hours with a standard deviation of 60 hours.

(a) Develop the null and alternative hypotheses.

(b) At α = 0.01, what is the rejection rule?

(c) What is the value of the test statistic χ2?

(d) Has the researcher collected sufficient data to reject H0?

(e) Based on your answer in part (d), what is our conclusion as it relates to this situation?

pls write out in paper.
Answer #1

-Aus Given that; mean = 150 sample size na 300 mean time - 158 S=60 standard denation = 70 (a). The sell hypothesis & alterna= 219.69) (d). 2 def if reject to 219167 2245.068 O No ..we fail to Reject to (C). Conclusion. There is not sufficient eviden

(OR)

A] NULL HYPOTHESIS H0: 0 = 70

ALTERNATIVE HYPOTHESIS Ha: σ< 70

b] REJECTION RULE: REJECT HO IF CALCULATED CHI SQUARE LESS THAN CRITICAL VALUE OF CHI SQUARE=245.068

c] Test staistic X = (n-1)*82104 = (300 – 1) * (60)/(70)2 = 256.29

d] Yes.  the researcher collected sufficient data to reject H0.

e] Conclusion: We have sufficient evidence to conclude that  the standard deviation is actually less than the 70hours at 0.01 level of significance.

NOTE:: I HOPE YOUR HAPPY WITH MY ANSWER....***PLEASE SUPPORT ME WITH YOUR RATING...

***PLEASE GIVE ME "LIKE"...ITS VERY IMPORTANT FOR ME NOW....PLEASE SUPPORT ME ....THANK YOU